How many litrs of 20% alcohol solution must be mixed with a 60% solution to obtain 40 litres of a 35% solution?
Answer by checkley71(8403)

.2x+.6(40-x)=.35*40
.2x+24-.6x=14
-.4x=14-24
-.4x=-10
x=-10/-.4
x=25 liters of 20% are needed.
proof.
.2*25+.6(40-25)=14
5+.6*15=14
5+9=14
14=14